OpenStudy (anonymous):

**will give medal** Find the x-intercept of 3x + 2y = 12 is (__, 0). A. 3 B. 2 C. 6 D. 4

OpenStudy (the_fizicx99):

3x + 2y = 12 -3x -3x 2y = -3x + 12 /2 /2 /2 y = -3/2x + 6 M or slope is -3/2 Y intercept is 6

OpenStudy (the_fizicx99):

Remember that in y = mx + b m is the slope b is the y intercept

OpenStudy (anonymous):

basically when is y = 0, you would be able to find the x-intercept so by plugging it in it leaves you with 3x = 12 and x = 4 D

OpenStudy (the_fizicx99):

Ohh x intercept. My bad >.<

OpenStudy (the_fizicx99):

Well to find the x intercept set y for 0 To find the y intercept set x for 0 So, 3x + 2y = 12 3x + 2(0) = 12 3x = 12 /3 /3 x = 4

OpenStudy (anonymous):

obvi that's what i did bro

OpenStudy (the_fizicx99):

shaddup ;~; , I misread/understood it thought it said y intercept

OpenStudy (anonymous):

lol it's cool man